what volume (ml) of 0.500M HCL is needed to neutralize 16.00mL of 0.200M Ba(OH)2 ?


1
Expert's answer
2021-12-28T14:16:01-0500

Reaction equation is :

Ba(OH)2 + 2HCl "\implies" BaCl2 + 2H2O

Mole ratio is:

1 : 2 for barium hydroxide and hydrochloric acid respectively.

Moles of Ba(OH)2 used is

0.2×16/1000 = 0.0032moles

Moles of acid used is

2 × 0.0032 = 0.0064moles

Volume used is

0.0064 × 1000/0.5 = 12.8 ml

Ans =12.8 ml